# Rethinking Text-to-SQL: Dynamic Multi-turn SQL Interaction for Real-world Database Exploration

### v1

@@ I Introduction
Structured Query Language (SQL) has become a cornerstone of data-driven applications across domains such as customer service analytics, electronic health record exploration, and financial risk monitoring [1, 2]. For domain experts with clearly defined analytical objectives, a single, well-crafted SQL query is often sufficient to obtain the desired information. However, in many real-world scenarios, users cannot fully articulate their intentions in a single turn due to incomplete prior knowledge, ambiguous goals, or evolving analytical requirements [3]. As a result, interactions with databases naturally unfold as multi-turn dialogues, where users iteratively refine query constraints, clarify ambiguities, and progressively explore the data space.
@@
Recent advances in LLMs have markedly improved performance on Text-to-SQL [4, 5, 6]. To assess these capabilities, several benchmarks have been introduced, including Spider 1 [7], Spider 2 [8], and BIRD [9]. Despite their impact, important gaps remain:
(1) Limited multi-turn coverage. As shown in Figure 1, most datasets and benchmarks focus on single-turn tasks or static multi-turn tasks [10] and fail to capture realistic conversational workflows in which users iteratively refine queries, correct mistakes, or add constraints across turns.
(2) Incomplete CRUD spectrum. Although MultiSQL [11] is an early effort toward covering the full CRUD spectrum, the majority of existing benchmarks still emphasize read-only $SELECT$ queries and provide little to no assessment of write operations, $INSERT/UPDATE/DELETE$, which frequently arise in practice.
@@
Figure 2: To assess the quality of synthesized realistic tool-use data, prior approaches typically relied on complex ad hoc verification procedures. In contrast, our Tool-Verifier is concise and efficient.
@@ II Related Work
Mapping natural language utterances into executable SQL queries has long been viewed as a promising way to democratize database access, freeing users from the burden of mastering schema intricacies and SQL syntax [12, 1]. Recent progress in LLMs [13, 14, 15, 16, 17] has substantially advanced this direction, driven by their powerful reasoning and cross-domain generalization abilities [18, 19, 20]. A number of recent efforts have sought to refine this paradigm by decomposing the problem and leveraging contextual reasoning [21, 22]. For instance, few-shot frameworks such as DIN-SQL [23] and DAIL-SQL [24] employ in-context demonstrations to separate schema linking from SQL generation, while DTS-SQL [25] enhance smaller-scale models through selective data curation. In parallel, agent-style systems that integrate thought, action, and feedback, like MAC-SQL [26], illustrating that iterative interaction with the environment can lead to notable performance gains [27]. Although these studies have collectively improved SQL synthesis accuracy, the majority of them remain confined to single-turn settings [28]. Consequently, their robustness and adaptability in multi-turn conversational scenarios have yet to be systematically explored.
@@
Benchmark development has been central to progress in Text-to-SQL research [29, 30, 31]. Early datasets such as ATIS [32], a flight-booking system dataset mapping natural‐language user queries about airline travel into structured queries, and GeoQuery [33], a U.S. geography question–answering dataset converting natural-language questions into formal queries, provided domain-specific testbeds that enabled early system design but lacked schema diversity and compositional depth. The introduction of SPIDER [7] fundamentally transformed the field by emphasizing cross-domain generalization to unseen databases, catalyzing advances in schema linking, compositional reasoning, and data augmentation. Building on this foundation, follow-up benchmarks like SPIDER 2.0 [8] and BIRD [9] introduced richer database schemas, paraphrased queries, and dynamic evaluation settings to approximate realistic database interaction. However, existing benchmarks remain largely confined to single-turn query formulation, assuming that a user’s intent can be fully captured in one utterance. In real-world scenarios, users often express goals progressively, refining or expanding their requests based on intermediate outcomes.
@@
In real-world applications, user queries are often ambiguous, incomplete, or evolve through conversation [34, 35]. Multi-turn Text-to-SQL research thus focuses on handling underspecified queries by leveraging clarification and context tracking. Early datasets such as CoSQL [10] extend the SPIDER [7] benchmark with dialogue-based turns to simulate this process. However, these benchmarks assume static and noise-free dialogue histories, neglecting that systems may initiate different clarification strategies [36]. More recent works investigating autonomous agents that maintain dynamic conversational states [34, 37], yet their methodologies have not been adapted to the Text-to-SQL context. Constructing an effective user simulator for this task remains non-trivial: it must balance database realism with controlled answer spaces and schema constraints. To bridge this gap, our study introduces a multi-turn benchmark featuring an user simulator, dynamic evaluation and real-world databases. This framework enables a systematic evaluation of reasoning-oriented models under realistic and uncertain Text-to-SQL conditions.
@@ IV-B Refined DB Task
Particularly, for each vanilla task, validation is first conducted by a Verifier Committee composed of multiple LLM-based validators. Each validator performs $n$ independent checks, and a task is considered verified only if all $n$ checks of two distinct validators, DeepSeek-r1 [18] and Qwen3-235B-A22B-2507 [19]. After this stage, we observed that even when golden actions could technically resolve the instruction, the instruction itself sometimes omitted necessary parameters due to limitations of the initial prompting. To address this issue, we introduce a data refinement stage, where parameters required by the golden actions are backfilled into the instruction. Following refinement, the updated tasks undergo a second round of committee validation to ensure that no additional hallucinations are introduced. After that, We then test the executability of golden actions in a mock runtime environment, discarding all tasks containing SQL syntax errors or execution failures.
@@ VI-A1 Hyperparameters
We conducted a systematic evaluation of a wide range of open-source and closed-source models, including GPT-4o [38], DeepSeek-V3 [39], Qwen2.5-Max, Qwen2.5-72B-Instruct [40], Llama-3.1-70B-Instruct [41], OmniSQL-32B [4], Qwen3-32B [19], Qwen3-Coder-30B-A3B-Instruct [19] and Gemini-2.5-Flash [42]. We set the maximum number of dialogue turns $\eta$ to 30, and fixed the temperature to 0.6, top_p to 0.95, and top_k to 20 for all tested models. Qwen2.5-72B-Instruct was employed as the model simulating the user.
@@ VI-A2 Metric
We adopt the Passˆk metric proposed in [34], this metric is defined as the probability that k i.i.d. solution samples for a given task are all correct, averaged across a distribution of tasks. Formally, it is calculated as follows:
@@ VI-A3 Implementation Details
For models with up to 70B parameters, all experiments are conducted on a single server equipped with eight NVIDIA H200 GPUs (each with 140 GB of memory). Using SGLang [43], we deploy the Qwen2.5-72B-Instruct model as the simulated user on four GPUs, while the remaining four GPUs are used to host the agent model under evaluation, also deployed via SGLang. For closed-source models and DeepSeek-r1, we conduct the evaluation through remote API calls.
@@ VI-C Consistency and Stability Analysis Across Multiple Trials
During dynamic multi-turn interactions between the agent model and the user model, we found that the agent often exhibits hallucination after generating an SQL query, namely by fabricating the query results on its own. According to the taxonomy of hallucinations defined in prior work [44], this phenomenon falls under Extrinsic Hallucination, as it involves generations that are inconsistent with the training data and cannot be verified by the given context. Such hallucinations typically arise when models attempt to fill knowledge gaps or produce unsupported content beyond the scope of the provided input, rather than misinterpreting the input itself. We identify hallucinations by checking whether each model’s dialogue trajectory output includes an SQL code block that is immediately followed by the special token <result>, which is used in the agent model’s system prompt to denote the SQL execution result returned from the environment. As illustrated in Figure 9, DeepSeek-V3 and OmniSQL-32B demonstrate the highest hallucination rates, at 44.28% and 26.15%, respectively. The hallucination rates of Llama3.1-70B-Instruct, Qwen3-Coder-30B-A3B-Instruct, Qwen2.5-Max, and Gemini-2.5-Flash are approximately 10%, whereas GPT-4o yields the lowest hallucination rate of 0.09%. We hypothesize that the high hallucination rates observed in DeepSeek-V3 and OmniSQL-32B stem from the models’ tendency to learn step-by-step problem-solving patterns during post-training. In this process, the models are trained to generate solutions in a procedural manner, often providing intermediate results at each step. Consequently, they internalize this paradigm. However, in the SQL-calling scenario, the execution results are produced externally by the environment rather than inferred by the model itself. When the model continues to follow its learned reasoning paradigm in this context, it tends to fabricate results, leading to hallucinations. We present an example in Figure 10. After receiving the user’s instruction, the model first performs identity verification. However, after generating a query to retrieve the user’s ID (BowlerID), it fabricates an incorrect result (BowlerID = 14), whereas the actual query output should be BowlerID = 8. This behavior demonstrates that the model produced a factually incorrect output due to hallucination.
@@ VI-E Analysis of Dialogue Turns and Erroneous SQL Invocation Turns
We analyze how different models perform when SQL execution fails across databases. For clarity and conciseness, we select three representative databases, entertainment, retail_world and car, and use GPT-4o, DeepSeek-V3, Qwen2.5-72B-Instruct, and Llama-3.1-70B-Instruct as examples to illustrate their behaviors on these databases. As shown in Figure 7, DeepSeek-V3 exhibits a pronounced long-tailed distribution. Its central 60% of dialogues span a wide range—approximately 2.7k–7.1k tokens on retail_world and 1.5k–4.9k tokens on car—while several extreme cases exceed 20k tokens. We attribute this phenomenon to the use of GRPO algorithm [45] during post-training, which likely encourages the model to generate longer and more exploratory responses, thereby substantially increasing dialogue length. In contrast, GPT-4o and Qwen2.5-72B-Instruct display compact, single-peaked distributions centered around 1.4k–1.9k tokens, reflecting stable clarify–execute behavior and strong schema grounding. Llama-3.1-70B-Instruct follows a similar pattern but with dialogue lengths typically around 1.3k–2.4k tokens. As shown in Figure 8, we observe that GPT-4o attains the highest fraction of zero-error turns on both car and retail_world; non-zero errors are rare, with only a small spike at three errors on car. Qwen2.5-72B-Instruct and Llama-3.1-70B-Instruct also concentrate at zero but exhibit a more visible single-error bar, consistent with minor, recoverable mismatches. DeepSeek-V3 likewise has many zero-error turns and few multi-error cases, yet its interactions are markedly longer. Compared with Qwen2.5-72B-Instruct, both DeepSeek-V3 and GPT-4o produce SQL invocations that fail less often due to syntax errors. However, their end-to-end task accuracy remains noticeably lower than Qwen2.5’s(see Table V). This gap suggests that on our benchmark, intent understanding, multi-turn planning, and schema-aware SQL synthesis—rather than surface-level SQL correctness—are the primary bottlenecks for DeepSeek-V3 and GPT-4o, indicating room for improvement in generating higher-quality, goal-satisfying SQL over extended dialogues. Extending to the entertainment domain, we observe that GPT-4o’s proportion of turns with zero SQL invocation errors is significantly lower than that of other models, while its proportions for one, two, and three errors remain comparable. Nevertheless, it maintains near-SOTA results on Passˆ1, achieving 80.43% accuracy on Short and 64.71% on Long samples (see Table III and  IV). This finding suggests that when the task inherently requires a larger number of SQL invocations, GPT-4o is able to balance invocation frequency and overall task performance more effectively, avoiding unnecessary SQL calls while maintaining strong end-to-end accuracy.

### v2

@@ I Introduction
Structured Query Language (SQL) has become a cornerstone of data-driven applications across domains such as customer service analytics, electronic health record exploration, and financial risk monitoring [katsogiannis2023survey, shi2025survey]. For domain experts with clearly defined analytical objectives, a single, well-crafted SQL query is often sufficient to obtain the desired information. However, in many real-world scenarios, users cannot fully articulate their intentions in a single turn due to incomplete prior knowledge, ambiguous goals, or evolving analytical requirements [zhang2025survey]. As a result, interactions with databases naturally unfold as multi-turn dialogues, where users iteratively refine query constraints, clarify ambiguities, and progressively explore the data space.
@@
Recent advances in LLMs have markedly improved performance on Text-to-SQL [li2025omnisql, ma2025sql, pourreza2025reasoning]. To assess these capabilities, several benchmarks have been introduced, including Spider 1 [yu2018spider], Spider 2 [lei2024spider], and BIRD [li2023can]. Despite their impact, important gaps remain:
(1) Limited multi-turn coverage. As shown in Figure 1, most datasets and benchmarks focus on single-turn tasks or static multi-turn tasks [yu2019cosql] and fail to capture realistic conversational workflows in which users iteratively refine queries, correct mistakes, or add constraints across turns.
(2) Incomplete CRUD spectrum. Although MultiSQL [li2024multisql] is an early effort toward covering the full CRUD spectrum, the majority of existing benchmarks still emphasize read-only $SELECT$ queries and provide little to no assessment of write operations, $INSERT/UPDATE/DELETE$, which frequently arise in practice.
@@
Figure 2: Task Generation Pipeline. Based on 13 databases from Spider 2 and BIRD, we designed a two-stage DynSQL-Bench task synthesis method. In the first stage, an LLM generates initial tasks. In the second stage, a filtering mechanism selects high-quality tasks.
@@ II Related Work
Mapping natural language utterances into executable SQL queries has long been viewed as a promising way to democratize database access, freeing users from the burden of mastering schema intricacies and SQL syntax [qin2022survey, katsogiannis2023survey]. Recent progress in LLMs [an2025unictokens, an2024mc, luo2024llm, lin2025perceiveanythingrecognizeexplain, lin2024draw] has substantially advanced this direction, driven by their powerful reasoning and cross-domain generalization abilities [guo2025deepseek, yang2025qwen3, chen2025learning]. A number of recent efforts have sought to refine this paradigm by decomposing the problem and leveraging contextual reasoning [sun2024beats, li2025swe]. For instance, few-shot frameworks such as DIN-SQL [pourreza2023din] and DAIL-SQL [gao2023text] employ in-context demonstrations to separate schema linking from SQL generation, while DTS-SQL [pourreza2024dts] enhance smaller-scale models through selective data curation. In parallel, agent-style systems that integrate thought, action, and feedback, like MAC-SQL [wang2023mac], illustrating that iterative interaction with the environment can lead to notable performance gains [yao2022react]. Although these studies have collectively improved SQL synthesis accuracy, the majority of them remain confined to single-turn settings [hong2025next]. Consequently, their robustness and adaptability in multi-turn conversational scenarios have yet to be systematically explored.
@@
Benchmark development has been central to progress in Text-to-SQL research [zhang2024benchmarking, mitsopoulou2025analysis, bhaskar2023benchmarking]. Early datasets such as ATIS [price1990evaluation], a flight-booking system dataset mapping natural‐language user queries about airline travel into structured queries, and GeoQuery [zelle1996learning], a U.S. geography question–answering dataset converting natural-language questions into formal queries, provided domain-specific testbeds that enabled early system design but lacked schema diversity and compositional depth. The introduction of SPIDER [yu2018spider] fundamentally transformed the field by emphasizing cross-domain generalization to unseen databases, catalyzing advances in schema linking, compositional reasoning, and data augmentation. Building on this foundation, follow-up benchmarks like SPIDER 2.0 [lei2024spider] and BIRD [li2023can] introduced richer database schemas, paraphrased queries, and dynamic evaluation settings to approximate realistic database interaction. However, existing benchmarks remain largely confined to single-turn query formulation, assuming that a user’s intent can be fully captured in one utterance. In real-world scenarios, users often express goals progressively, refining or expanding their requests based on intermediate outcomes.
@@
In real-world applications, user queries are often ambiguous, incomplete, or evolve through conversation [yao2024tau, barres2025tau]. Multi-turn Text-to-SQL research thus focuses on handling underspecified queries by leveraging clarification and context tracking. Early datasets such as CoSQL [yu2019cosql] extend the SPIDER [yu2018spider] benchmark with dialogue-based turns to simulate this process. However, these benchmarks assume static and noise-free dialogue histories, neglecting that systems may initiate different clarification strategies [yi2024survey]. More recent works investigating autonomous agents that maintain dynamic conversational states [yao2024tau, li2025beyond], yet their methodologies have not been adapted to the Text-to-SQL context. Constructing an effective user simulator for this task remains non-trivial: it must balance database realism with controlled answer spaces and schema constraints. To bridge this gap, our study introduces a multi-turn benchmark featuring an user simulator, dynamic evaluation and real-world databases. This framework enables a systematic evaluation of reasoning-oriented models under realistic and uncertain Text-to-SQL conditions.
@@ IV-B Refined DB Task
Particularly, for each vanilla task, validation is first conducted by a Verifier Committee composed of multiple LLM-based validators. Each validator performs $n$ independent checks, and a task is considered verified only if all $n$ checks of two distinct validators, DeepSeek-r1 [guo2025deepseek] and Qwen3-235B-A22B-2507 [yang2025qwen3]. After this stage, we observed that even when golden actions could technically resolve the instruction, the instruction itself sometimes omitted necessary parameters due to limitations of the initial prompting. To address this issue, we introduce a data refinement stage, where parameters required by the golden actions are backfilled into the instruction. Following refinement, the updated tasks undergo a second round of committee validation to ensure that no additional hallucinations are introduced. After that, We then test the executability of golden actions in a mock runtime environment, discarding all tasks containing SQL syntax errors or execution failures.
@@ VI-A1 Hyperparameters
We conducted a systematic evaluation of a wide range of open-source and closed-source models, including GPT-4o [hurst2024gpt], DeepSeek-V3 [liu2024deepseek], Qwen2.5-Max, Qwen2.5-72B-Instruct [qwen2.5], Llama-3.1-70B-Instruct [vavekanand2024llama], OmniSQL-32B [li2025omnisql], Qwen3-32B [yang2025qwen3], Qwen3-Coder-30B-A3B-Instruct [yang2025qwen3] and Gemini-2.5-Flash [comanici2025gemini]. We set the maximum number of dialogue turns $\eta$ to 30, and fixed the temperature to 0.6, top_p to 0.95, and top_k to 20 for all tested models. Qwen2.5-72B-Instruct was employed as the model simulating the user.
@@ VI-A2 Metric
We adopt the Passˆk metric proposed in [yao2024tau], this metric is defined as the probability that k i.i.d. solution samples for a given task are all correct, averaged across a distribution of tasks. Formally, it is calculated as follows:
@@ VI-A3 Implementation Details
For models with up to 70B parameters, all experiments are conducted on a single server equipped with eight NVIDIA H200 GPUs (each with 140 GB of memory). Using SGLang [zheng2024sglang], we deploy the Qwen2.5-72B-Instruct model as the simulated user on four GPUs, while the remaining four GPUs are used to host the agent model under evaluation, also deployed via SGLang. For closed-source models and DeepSeek-r1, we conduct the evaluation through remote API calls.
@@ VI-C Consistency and Stability Analysis Across Multiple Trials
During dynamic multi-turn interactions between the agent model and the user model, we found that the agent often exhibits hallucination after generating an SQL query, namely by fabricating the query results on its own. According to the taxonomy of hallucinations defined in prior work [bang2025hallulens], this phenomenon falls under Extrinsic Hallucination, as it involves generations that are inconsistent with the training data and cannot be verified by the given context. Such hallucinations typically arise when models attempt to fill knowledge gaps or produce unsupported content beyond the scope of the provided input, rather than misinterpreting the input itself. We identify hallucinations by checking whether each model’s dialogue trajectory output includes an SQL code block that is immediately followed by the special token <result>, which is used in the agent model’s system prompt to denote the SQL execution result returned from the environment. As illustrated in Figure 9, DeepSeek-V3 and OmniSQL-32B demonstrate the highest hallucination rates, at 44.28% and 26.15%, respectively. The hallucination rates of Llama3.1-70B-Instruct, Qwen3-Coder-30B-A3B-Instruct, Qwen2.5-Max, and Gemini-2.5-Flash are approximately 10%, whereas GPT-4o yields the lowest hallucination rate of 0.09%. We hypothesize that the high hallucination rates observed in DeepSeek-V3 and OmniSQL-32B stem from the models’ tendency to learn step-by-step problem-solving patterns during post-training. In this process, the models are trained to generate solutions in a procedural manner, often providing intermediate results at each step. Consequently, they internalize this paradigm. However, in the SQL-calling scenario, the execution results are produced externally by the environment rather than inferred by the model itself. When the model continues to follow its learned reasoning paradigm in this context, it tends to fabricate results, leading to hallucinations. We present an example in Figure 10. After receiving the user’s instruction, the model first performs identity verification. However, after generating a query to retrieve the user’s ID (BowlerID), it fabricates an incorrect result (BowlerID = 14), whereas the actual query output should be BowlerID = 8. This behavior demonstrates that the model produced a factually incorrect output due to hallucination.
@@ VI-E Analysis of Dialogue Turns and Erroneous SQL Invocation Turns
We analyze how different models perform when SQL execution fails across databases. For clarity and conciseness, we select three representative databases, entertainment, retail_world and car, and use GPT-4o, DeepSeek-V3, Qwen2.5-72B-Instruct, and Llama-3.1-70B-Instruct as examples to illustrate their behaviors on these databases. As shown in Figure 7, DeepSeek-V3 exhibits a pronounced long-tailed distribution. Its central 60% of dialogues span a wide range—approximately 2.7k–7.1k tokens on retail_world and 1.5k–4.9k tokens on car—while several extreme cases exceed 20k tokens. We attribute this phenomenon to the use of GRPO algorithm [shao2024deepseekmath] during post-training, which likely encourages the model to generate longer and more exploratory responses, thereby substantially increasing dialogue length. In contrast, GPT-4o and Qwen2.5-72B-Instruct display compact, single-peaked distributions centered around 1.4k–1.9k tokens, reflecting stable clarify–execute behavior and strong schema grounding. Llama-3.1-70B-Instruct follows a similar pattern but with dialogue lengths typically around 1.3k–2.4k tokens. As shown in Figure 8, we observe that GPT-4o attains the highest fraction of zero-error turns on both car and retail_world; non-zero errors are rare, with only a small spike at three errors on car. Qwen2.5-72B-Instruct and Llama-3.1-70B-Instruct also concentrate at zero but exhibit a more visible single-error bar, consistent with minor, recoverable mismatches. DeepSeek-V3 likewise has many zero-error turns and few multi-error cases, yet its interactions are markedly longer. Compared with Qwen2.5-72B-Instruct, both DeepSeek-V3 and GPT-4o produce SQL invocations that fail less often due to syntax errors. However, their end-to-end task accuracy remains noticeably lower than Qwen2.5’s(see Table V). This gap suggests that on our benchmark, intent understanding, multi-turn planning, and schema-aware SQL synthesis—rather than surface-level SQL correctness—are the primary bottlenecks for DeepSeek-V3 and GPT-4o, indicating room for improvement in generating higher-quality, goal-satisfying SQL over extended dialogues. Extending to the entertainment domain, we observe that GPT-4o’s proportion of turns with zero SQL invocation errors is significantly lower than that of other models, while its proportions for one, two, and three errors remain comparable. Nevertheless, it maintains near-SOTA results on Passˆ1, achieving 80.43% accuracy on Short and 64.71% on Long samples (see Table III and  IV). This finding suggests that when the task inherently requires a larger number of SQL invocations, GPT-4o is able to balance invocation frequency and overall task performance more effectively, avoiding unnecessary SQL calls while maintaining strong end-to-end accuracy.